OpenStudy (anonymous):

If a car is travelling 80 miles per hour on a slippery road, then its stopping distance D in feet can be calculated by D=(6400/80)*(1/x), where x is teh coefficient of friction between the tires and road and 0

OpenStudy (dumbcow):

D = 80/x you are just reducing x=.2 -> D = 80*5=400 x=.8 -> D=80*(5/4) = 100

OpenStudy (anonymous):

what about if teh car is moving at 80 miles per hour the stopping distance on dry pavement is __ of the stopping distance of an icy road?? (type as an integer or simplified fraction)

OpenStudy (anonymous):

I really appreciate your help

OpenStudy (dumbcow):

i already answered that....we are already assuming car is going 80 mph its already factored into equation for D